### v3

# Quantum Delocalization Enables Water Dissociation on Ru(0001)

Abstract
We revisit the long-standing question of whether water molecules dissociate on the Ru(0001) surface through nanosecond-scale path-integral molecular dynamics simulations on a sizable supercell. This is made possible through the development of an efficient and reliable machine-learning potential with near first-principles accuracy, overcoming the limitations of previous ab initio studies. We show that the quantum delocalization associated with nuclear quantum effects enables rapid and frequent proton transfers between water molecules, thereby facilitating the water dissociation on Ru(0001). This work provides the direct theoretical evidence of water dissociation on Ru(0001), resolving the enduring issue in surface sciences and offering crucial atomistic insights into water-metal interfaces.
The interaction of water with solid surfaces plays an important role in various scientific and technical fields, including ice nucleation, catalysis, lubrication, and corrosion Mu et al. (2017); Knopf and Alpert (2023); Meng and Wang (2023). The subtle interplay between the water-solid interface and water-water hydrogen bonding interactions results in complex structural motifs of water on surfaces, prompting extensive experimental and theoretical investigations Hodgson and Haq (2009); Carrasco et al. (2012); Maier and Salmeron (2015); Björneholm et al. (2016); Shimizu et al. (2018); Zhou and Huang (2021); Tian et al. (2022); Gao and Meng (2024). Among these studies, the water overlayer on the Ru(0001) surface is notably one of the most debated systems.
Initially, it was commonly believed that the first-layer water molecules on Ru(0001) would arrange themselves into a bulk-ice-like bilayer Thiel et al. (1982); Doering and Madey (1982); Thiel et al. (1984); Pirug et al. (1991). However, this notion was challenged by the low-energy electron diffraction (LEED) analysis by Held and Menzel Held and Menzel (1994); Puisto et al. (2003), which revealed nearly coplanar O atoms with a vertical distance of just 0.14 $\AA$ Puisto et al. (2003), significantly smaller than that of the ice-like bilayer. Moreover, density functional theory (DFT) calculations employing commonly used generalized gradient approximation functionals predicted that the ice-like bilayer did not even wet the surface Feibelman (2002). To resolve this discrepancy, Feibelman proposed a partially dissociated overlayer model Feibelman (2002), which not only reproduced the experimental observation of almost coplanar O atoms, but also demonstrated superior energetic stability compared to the H-up or H-down bilayers as well as the sublimation energy of ice-Ih Feibelman (2002).
Feibelman’s study Feibelman (2002) raised an even today debated question—whether the water molecules within the overlayer dissociate on Ru(0001)? Experimental evidence from sum-frequency generation vibrational spectroscopy Denzler et al. (2003a) and reflection adsorption infrared spectroscopy Clay et al. (2004); Haq et al. (2006) indicated that the water remained intact on Ru(0001) at low temperatures. However, temperature programmed desorption measurements observed residual atomic hydrogen on the surface after water desorption, implying the dissociative adsorption Denzler et al. (2003b). The situation became even more puzzling by conflicting x-ray photoelectron spectroscopy (XPS) studies Andersson et al. (2004); Faradzhev et al. (2005); Tatarkhanov et al. (2008); Weissenrieder et al. (2004). Andersson et al. attributed the flat water-hydroxyl overlayer observed by LEED to electron beam damage and claimed that the water dissociation was a thermal or electron stimulated process Andersson et al. (2004), which was later supported by two independent XPS studies Faradzhev et al. (2005); Tatarkhanov et al. (2008). In contrast, Weissenrieder et al. provided experimental XPS evidence for a partially dissociated water bilayer on Ru(0001) at temperatures as low as 105 K using beam-damage-free low photon flux Weissenrieder et al. (2004). Besides XPS, scanning tunneling microscopy (STM) also unveiled perplexing results. Tatarkhanov et al. Tatarkhanov et al. (2009) and Maier et al. Maier et al. (2012, 2014) demonstrated through STM that at temperatures below 130 K, the deposited water remained intact and tended to form stripes of hexagons. However, a more recent STM investigation by Schilling and Behm Schilling and Behm (2018) indicated that partial water dissociation could possibly occur already during slow adsorption at 120 K.
Figure 1: Top and side views of (a) H-down, (b) H-up, (c) extended chains, (d) $\sqrt{3}{\times}\sqrt{3}$ half-dissociated Weissenrieder et al. (2004), and (e) $3\sqrt{3}\times 3$ half-dissociated Feibelman (2002) water overlayer models. The black lines indicate the unit cell. (f) The DFT and MTP predicted adsorption energies of H2O on Ru(0001) for the 43 considered configurations at various coverages (see Supplemental Material Fig. S2 SM ). The dashed line denotes the sublimation energy of ice-Ih.
The conflicting experimental results also stimulated extensive DFT calculations Michaelides et al. (2003a, 2004); Meng et al. (2004, 2005); Michaelides et al. (2003b); Feibelman (2005); Gallagher et al. (2009); Carrasco et al. (2011); Herron et al. (2013). Consistent with Feibelman’s predictions, Michaelides et al. Michaelides et al. (2003a, 2004) confirmed that the partially dissociated overlayer was thermodynamically more stable than any identified water bilayers. Nevertheless, the predicted dissociation energy barrier was higher than the mean adsorption energy of the water bilayers Meng et al. (2004, 2005), indicating the favorable desorption of water over dissociation. To address the problems associated with the coplanar overlayer observed by LEED and the underestimated work function change upon adsorption for the half-dissociated model, a metastable structure with mixed intact H-up and H-down molecules Meng et al. (2005), and an extended chains model containing chains of flat-lying and chains of H-down molecules Haq et al. (2006); Gallagher et al. (2009) were suggested.
Despite comprehensive DFT studies, the ground-state structure of water overlayers on Ru(0001) remains inconclusive due to the vast configuration space originating from the complex hydrogen bonding networks and water-metal interactions. Additionally, whether the water overlayers dissociate on Ru(0001) remains elusive, given the high predicted dissociation energy barrier (0.5 eV Michaelides et al. (2003a) and 0.62 eV Meng et al. (2005)). Aiming to address these questions, we developed an accurate, efficient, and robust machine-learning potential, which not only facilitates the accurate exploration of the potential energy surface (PES) landscape for water adsorption on Ru(0001) at various coverages, but also allows for long-time scale path-integral molecular dynamics (PIMD) simulations. Our work provides direct and compelling evidence of water dissociation from an intact overlayer on Ru(0001), which is driven by nuclear quantum effects (NQEs).
We begin by constructing and validating our moment tensor potential (MTP) Shapeev (2016). Following the scheme outlined in Refs. Liu et al. (2023, 2024), the training dataset was initially acquired through an on-the-fly active learning procedure Jinnouchi et al. (2019a, b) implemented in the Vienna ab initio simulation package (VASP) Kresse and Hafner (1993); Kresse and Furthmüller (1996), and subsequently expanded iteratively using the active learning method implemented in the MLIP package Novikov et al. (2021). First-principles calculations were performed using VASP Kresse and Hafner (1993); Kresse and Furthmüller (1996) employing the revised Perdew-Burke-Ernzerhof (RPBE) functional for its improved adsorption energetics Hammer et al. (1999). Van der Waals (vdW) interactions were included using the D3 method of Grimme et al. Grimme et al. (2010) with zero damping. We found that RPBE+D3 yielded similar Ru-O and O-O vertical distances and relative water adsorption energies when compared to SCAN Sun et al. (2015), PBE+D3 Perdew et al. (1996); Grimme et al. (2010), and optB88-vdW Klimeš et al. (2009) functionals (Supplemental Material Table S1 and Fig. S1 SM ). The training dataset contained 6,813 structures including bulk ice-Ih, bulk Ru, and the clean Ru(0001) surface, as well as intact and dissociated water overlayers at various coverages up to 2/3 monolayer (ML) (Supplemental Material Table S2 SM ). To enhance the efficiency, the MTP was fitted with the optimized basis sets by refining the contraction process of moment tensors using our in-house code (IMR-MLP) Wang et al. . The MTP was validated on a test dataset of 600 structures including bulk Ru, the clean Ru(0001) surface, intact and half-dissociated water overlayers at 2/3 ML (Supplemental Material Table S3 SM ). These test structures were randomly sampled from MD simulations. The MTP demonstrates a high accuracy, with validation root-mean-square errors of 1.02 meV/atom for energies and 59 $\AA$ for forces (Supplemental Material Fig. S3 SM ). More computational details can be found in the Supplemental Material SM .
Figure 2: Time evolution of the number of OH (CNO=1), H2O (CNO=2), and H3O (CNO=3) species in (a) classical MD and (b) PIMD simulations at 300 K. CNO represents the coordination number of the oxygen atom. (c) Time evolution of occurrences when either of the two hydrogen atoms in an H2O molecule differs from the initial configuration at 0 ns. (d) Mean square displacements (MSD) of the O and H atoms. (e) Time evolution of the vertical distances of the H and O atoms [specified in (f)] in the PIMD simulation. (f)-(k) Snapshots from the PIMD trajectory. The tracked O and H atoms as well as the H3O and OH species are highlighted. The cyan arrows depict the proton transfer process.
The developed MTP allows us to efficiently explore the PES of water overlayers at different coverages. As depicted in Fig. 1, for the considered 43 configurations including intact and dissociated water overlayers, the MTP accurately predicts their adsorption energies, showing remarkable agreement with DFT results. Particularly, we focus on the extensively studied 2/3 ML overlayers. It is evident that the H-down and H-up models exhibit comparable adsorption energies. However, both configurations would not wet the Ru(0001) surface, since their adsorption energies are lower than the sublimation energy of ice-Ih [see Fig. 1(f) and Supplemental Material Fig. S1(a) SM ]. The most stable intact overlayer identified at 2/3 ML is the extended chains model with a unit cell of $\sqrt{3}{\times}2\sqrt{3}$, where the O atoms in the flat-lying chains are positioned atop Ru, whereas the O atoms in the H-down chains occupy the bridge sites [see Fig. 1(c)]. Note that this extended chains model is more stable by 0.04 eV/H2O than the one originally proposed by Hodgson et al. (i.e., cfg32) Haq et al. (2006); Gallagher et al. (2009), where all the O atoms are situated atop Ru. Generally, the dissociated overlayers (configurations 41-43 Feibelman (2002); Weissenrieder et al. (2004)) exhibit coplanar O atoms and are more stable compared to the intact ones, consistent with previous calculations Feibelman (2002); Michaelides et al. (2003a, 2004); Meng et al. (2004, 2005).
We note that the overlayer configurations examined above are constrained to ordered phases with small unit cells, without fully incorporating the disorder effects. Experiments also face challenges in identifying hydrogen disorder due to its light mass Held et al. (2001). To examine the disorder effects, we performed simulated annealing (SA) MD simulations on large supercells using the MTP and the LAMMPS code Thompson et al. (2022). The most favorable intact (cfg40) and dissociative (cfg43) configurations were selected as initial structures. Two large supercells (i.e., $6{\times}3\sqrt{3}$ and $12{\times}6\sqrt{3}$) were employed to account for varying degrees of disorder. For each configuration, 90 independent SA runs were conducted, followed by structural relaxations. Interestingly, for the dissociative case, the lowest-energy configuration identified by SA exhibits only marginally higher stability (0.01 eV/H2O energy gain) in comparison to the initial ordered phase (Supplemental Material Fig. S4 SM ). The slightly higher stability of cfg43 compared to cfg42 is also such a manifestation of disorder effects (Fig. 1). The disordered configurations accommodated within large supercells primarily exhibit disorder in hydrogen orientation, aimed at maximizing hydrogen bonding. However, for the intact case, we were unable to find a disordered configuration more favorable than the $\sqrt{3}{\times}2\sqrt{3}$ extended chains model within the 90 SA runs (Supplemental Material Fig. S4 SM ). This indicates a flat PES with numerous local minima, in agreement with Hodgson’s findings Gallagher et al. (2009). In fact, the extended chains configuration is a model that already accounts for partial disorder in the O height and proton orientation Gallagher et al. (2009). For both the dissociative and intact cases, a supercell beyond the $6{\times}3\sqrt{3}$ unit cell almost shows no energy gain from disorder.
Having identified the most stable intact and dissociative overlayers, we now turn to tackling the enduring question of whether water overlayers dissociate on Ru(0001). First, we computed the dissociation energy barrier, yielding 0.53 eV for dissociating half of the water molecules, aligning with previous predictions Michaelides et al. (2003a); Meng et al. (2005). The energy barrier is found to be higher ($>$0.6 eV) for dissociating 1/8 of the water molecules (Supplemental Material Fig. S5 SM ). Based on a rough estimation from transition state theory Truhlar et al. (1996), a dissociation energy barrier of this magnitude would necessitate a time scale exceeding microseconds, far beyond the accessible time-scale of MD simulations. Indeed, we did not observe the dissociation throughout a one-microsecond classical MD simulation [Fig. 2(a)]. Nevertheless, classical MD simulations do not take into account the NQEs, which have been shown to be significant in hydrogen-related systems Marx and Parrinello (1996); Benoit et al. (1998); Marx et al. (1999); Kumagai et al. (2008); Morrone et al. (2009); Li et al. (2010, 2011); Chen et al. (2013); Feng et al. (2015); Guo et al. (2016); Ceriotti et al. (2016); Guo et al. (2017); Chen et al. (2018); Markland and Ceriotti (2018); Fang et al. (2019); Cheng et al. (2019); Lan et al. (2020); Bore and Paesani (2023); Flór et al. (2024). In particular, Li et al. Li et al. (2010) showcased the pronounced NQEs for water overlayers on transition metal surfaces through ab initio PIMD simulations. Despite their high accuracy, ab initio methods face limitations concerning the accessible time and length scales, hindering the direct observation of water dissociations.
Figure 3: Probability distributions of (a) O-H and (b) O-O distances at 300 K. (c) Probability distribution of the proton transfer coordinate $|\delta|=|R_{\rm O_{a}H}-R_{\rm O_{b}H}|$, where $R_{\rm O_{a}H}$ denotes the distance between the Oa atom and the H atom [see the insert in (d)]. $|\delta|$=0 indicates that the shared H atom is exactly positioned between two O atoms. (d) Free energy profile for the protons along the intermolecular axes.
Thanks to our accurate and efficient machine-learning potential, we were able to perform PIMD simulations on large supercells for an extended timescale, while maintaining ab initio accuracy. In the PIMD simulations, 16 beads were employed to sample the imaginary time path-integral and the statistical analyses were derived by averaging these 16 beads. We employed a $6\times 4\sqrt{3}$ supercell with 32 H2O molecules based on the most stable intact overlayer configuration (i.e., extended chains model) as the initial structure. Throughout a 200-nanosecond PIMD simulation, three abrupt changes in the number of OH, H2O, and H3O species are observed [see Fig. 2(b)], indicating three instances of water dissociation. To corroborate our findings, we conducted additional independent PIMD and classical MD simulations. Water dissociation was consistently reproduced in all PIMD simulations but was absent in classical MD simulations within the considered timescale (Supplemental Material Fig. S7 SM ). Furthermore, PIMD simulations of D2O adsorption on Ru(0001), conducted under identical computational settings as H2O adsorption, revealed no dissociation. This result underscores the pronounced kinetic isotope effect and aligns with experimental observations Clay et al. (2004). Collectively, these results demonstrate that the NQEs facilitate the water dissociation on Ru(0001).
It is worth noting that such NQEs cannot be simply interpreted as the commonly assumed thermalization effect in classical MD simulations at elevated temperatures. This is because, even when the temperature is raised to 350 K, water dissociation does not occur within the same time frame in the classical MD simulation (Supplemental Material Fig. S7 SM ). Importantly, the NQEs result in a significant quantum delocalization of protons. This is evidenced by the broadened first peak in the probability distributions of O-H distances [Fig. 3(a)] as well as the enhanced probability distributions of protons within the central region of two adjacent oxygen atoms [Fig. 3(c)] observed in PIMD simulations, as compared to classical MD simulations. Moreover, the quantum delocalization of the shared protons between the two oxygen atoms slightly shortens the O-O distances [Fig. 3(b)], strengthening the hydrogen bonding. Our results agree well with the findings of Li et al. Li et al. (2010). Furthermore, we computed the free energy profile for the protons along the intermolecular axes through $\Delta F(|\delta|)=-k_{B}T\ln[P(|\delta|)]$ following Ref. Li et al. (2010), where $P(|\delta|)$ represents the probability distribution of $|\delta|=|R_{\rm O_{a}H}-R_{\rm O_{b}H}|$ [see Fig. 3(c)] and $k_{B}$ is the Boltzmann constant. One can see from Fig. 3(d) that the free energy barrier for proton transfer is just 59 meV in the PIMD simulation, a value notably lower than that observed in the classical MD simulation ($\sim$103 meV).
The small free energy barrier in PIMD simulations enables rapid and frequent proton transfers among water molecules [Figs. 2(c)-(d)]. Moreover, the emergence of water dissociation would enhance the rate of proton transfers [Fig. 2(c)]. Conversely, in classical MD simulations, the protons almost do not undergo transfers due to the large free energy barrier [Fig. 2(c)]. Specifically, in classical MD, the two H atoms and an O atom in an H2O molecule consistently exhibit identical motions, whereas in PIMD, they follow distinct motion pathways due to the proton transfers [Fig. 2(d)]. Interestingly, we observe that while the dissociated H atoms diffuse most rapidly with a diffusion coefficient of 5.58×$\times 10^{-5}$ cm2/s, the diffusion of H and O atoms within the water-hydroxyl overlayer is comparatively slower than within the intact overlayer (Supplemental Material Fig. S8 SM ). This might be related to the inhibited proton transfer by hydroxide ions Chen et al. (2018).
To unveil the details of proton transfers and water dissociation, we analyzed the PIMD trajectory. Note that although PIMD does not provide a real-time picture of the trajectories, we still believe that the mechanisms we observe would prevail in a centroid molecular dynamics and that the observed mechanisms are highly relevant to water dissociation. Initially, ephemeral hydroxyls exist due to the quantum delocalization associated with the NQEs, as evidenced by the fluctuation between the number of hydroxyls and intact H2O (Supplemental Material Fig. S9 SM ). At 3.8975 ns, the formation of an H3O and OH pair commences due to the proton transfer [Figs. 2(f)-(g)], and the pair would typically last for $\sim$38 ps (Supplemental Material Fig. S9 SM ). At 3.9355 ns, the first instance of water dissociation emerges. Interestingly, we observe that the dissociation does not directly originate from H3O, but instead undergoes an indirect process. Specifically, one of the protons in H3O is transferred to the neighboring H2O, leading to the simultaneous dissociation of a hydrogen atom from the latter [Figs. 2(h)-(i)]. The dissociated H is observed to preferentially occupy the hollow fcc sites rather than the top sites on Ru(0001). This observation is consistent with our DFT calculations, when the dissociation fraction is low [Supplemental Material Fig. S5(l) SM ]. Moreover, we find that once dissociated, the dissociated H decreases its vertical height [Fig. 2(e)], diffuses on the surface, but is rarely involved in proton transfers. The existence of the dissociated hydroxyl increases the frequency of proton transfers [Fig. 2(c)]. At 29.1005 ns, the second occurrence of water dissociation takes place, following a procedure akin to the first instance [Figs. 2(j)-(k) and Supplemental Material Fig. S10 SM ]. At 179.9685 ns, a third water molecule dissociates [Fig. 2(b)]. Notably, this third dissociation event does not involve the formation of H3O (Supplemental Material Fig. S11 SM ). At this point, the system contains 29 intact water molecules, 3 hydroxyls, and 3 dissociated hydrogen atoms. Over extended time scales, a greater number of H2O would likely dissociate. Note that within the simulated time frame, no water desorption was observed due to the higher energy barrier for desorption compared to dissociation (Supplemental Material Figs. S6 and S5 SM ). This aligns with experimental findings that H2O dissociates prior to desorption Clay et al. (2004). However, the large computational cost of PIMD kept us from much longer simulations.
In summary, we have developed an efficient and robust machine-learning potential to accurately capture the intricate interactions between water overlayers and the Ru(0001) surface through active learning and basis sets optimization. This enables us to identify the most stable overlayers for both intact and dissociative adsorption and facilitate the execution of PIMD simulations with near first-principles accuracy on a time and length scale that is inaccessible through ab initio methods. This capability is crucial, as it allows us to directly observe water dissociation on Ru(0001) through a nanosecond-scale PIMD simulation, a remarkable achievement that is unattainable with previous ab initio studies. We clearly demonstrate that the water dissociation is driven by the quantum delocalization originating from NQEs. These are essential to facilitate rapid and frequent proton transfers. Furthermore, the formation of H and OH pairs appears to occur via long-range density fluctuations, with an H and OH pair forming quasi-instantly over a distance of up to 5 connected H2O molecules. This is facilitated by a combination of the Grotthuss mechanism and quantum fluctuations. To our knowledge, such a long-range mechanism has not been observed before. Our study implies that quantum fluctuations can dramatically change the barriers to water dissociation, but also lead to new dissociation pathways. This has far-reaching implications not only for surface science, electrolysis and fuel cells, but also for biology, where water is ubiquitous. Clearly, our study suggests that quantum fluctuations must be taken into account whenever water dissociation occurs. The implications for the now so important field of hydrogen production are huge, as the barriers appear to be cut in half by quantum fluctuations.
This work is supported by the National Natural Science Foundation of China (Grants No. 52422112, No. 52188101, and No. 52201030), the Strategic Priority Research Program of the Chinese Academy of Sciences (XDA041040402), the Liaoning Province Science and Technology Major Project (2024JH1/11700032, 2023021207-JH26/103 and RC230958), the National Key R&D Program of China 2021YFB3501503, and the Special Projects of the Central Government in Guidance of Local Science and Technology Development (2024010859-JH6/1006).
Data availability—The data that support the findings of this article are openly available Dat .